OpenStudy (anonymous):

What is the measure of one exterior angle of a regular hexagon? 60 72 90

OpenStudy (anonymous):

@SpooksDF @Gralwyn

OpenStudy (anonymous):

the measure of 1 interior angle is 108, so the exterior must by 360 -108

OpenStudy (anonymous):

(n-2)180 = 180n

OpenStudy (anonymous):

so 252

OpenStudy (anonymous):

wait, a hexagon?

OpenStudy (anonymous):

sorry, I was wrong

OpenStudy (anonymous):

yep,

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

that would be (6-2)30 = 120 per each interior angle so 240 for an exterior angle

OpenStudy (anonymous):

(n-2) * 180/n

OpenStudy (anonymous):

4*180/6

OpenStudy (anonymous):

So it would be 120

OpenStudy (anonymous):

but i dont have the choice 120

OpenStudy (anonymous):

@SpooksDF @Gralwyn

OpenStudy (anonymous):

the choices are wrong?

OpenStudy (anonymous):

(n-2)(180)/n = 4(180)/6 = 120 360 - 120 = 240

OpenStudy (anonymous):

Yeah, The choices just suck

OpenStudy (anonymous):

thnx:)